# Supervised Learning-enhanced Multi-Group Actor Critic for Live-stream Recommendation

### v1

@@ Abstract.
Reinforcement Learning (RL) has been widely applied in recommendation systems to capture users’ long-term engagement, thereby improving dwelling time and enhancing user retention. In the context of a short video & live-stream mixed recommendation scenario, the live-stream recommendation system (RS) decides whether to inject at most one live-stream into the video feed for each user request. To maximize long-term user engagement, it is crucial to determine an optimal live-stream injection policy for accurate live-stream allocation. However, traditional RL algorithms often face divergence and instability problems, and these issues are even more pronounced in our scenario. To address these challenges, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm (SL-MGAC). Specifically, we introduce a supervised learning-enhanced actor-critic framework that incorporates variance reduction techniques, where multi-task reward learning helps restrict bootstrapping error accumulation during critic learning. Additionally, we design a multi-group state decomposition module for both actor and critic networks to reduce prediction variance and improve model stability. Empirically, we evaluate the SL-MGAC algorithm using offline policy evaluation (OPE) and online A/B testing. Experimental results demonstrate that the proposed method not only outperforms baseline methods but also exhibits enhanced stability in online recommendation scenarios.
@@
Figure 1. Structure of a short video & live-stream mixed recommendation system(RS). The decision making of SL-MGAC takes place in the final stage of live-stream RS.
@@ 1. Introduction
Deep reinforcement learning has shown great potential in learning policies to maximize long-term rewards across various research domains, such as computer vision, robotics, natural language processing, gaming, and recommendation systems (Afsar et al., 2022; Li, 2017). In the context of recommendation systems, reinforcement learning is applied to optimize long-term user engagement (Zou et al., 2019) and improve user retention (Cai et al., 2023). Numerous RL applications have been proposed for real-world recommendation systems, including slate recommendation (Ie et al., 2019; Deffayet et al., 2023; Liu et al., 2023), personalized search ranking (Miao et al., 2021), and advertisement allocation (Liao et al., 2022).
@@
We consider a challenging sequential decision-making scenario in a short video & live-stream blended recommendation system, as shown in Fig. 1. The system consists of three components: a live-stream recommendation system, a short video recommendation system, and a blending server. For each user request at timestamp $t$, the live-stream recommendation system decides whether to inject the recommended live-stream into the video feed, while the short video recommendation system suggests $B$ (with $B<10$) videos. The blending server then mixes and rearranges the single live-stream with the $B$ videos to form the final recommendation list. Our objective is to find a personalized optimal live-stream injection policy that maximizes users’ long-term engagement with live-streams, which can be modeled as an infinite request-level Markov Decision Process (MDP) (Puterman, 1990).
@@
However, the intrinsic issues of divergence and instability associated with traditional reinforcement learning (RL) algorithms (François-Lavet et al., 2015; Fujimoto et al., 2018; Kumar et al., 2019; Dasagi et al., 2019) are significantly exacerbated in this challenging scenario. A primary possible reason for this is the drastic fluctuations in both the live-stream supply scale and user interaction behaviors over time. As shown in Fig. 2a, the curves of live-stream room count and live-stream viewer count vary significantly on a daily basis, making it difficult for the RL algorithm to learn optimal policies. Furthermore, as illustrated in Fig. 2b, the distributions of user interaction labels (e.g., live-stream and short video watch time) are highly noisy and exhibit large variance. RL models often struggle to learn effective policies from data that exhibit such high variance across both time and user scales. In practice, we observe that RL models frequently encounter issues of policy deterioration or model collapse (Dohare et al., 2023).
Figure 2. Data distributions of online live-stream room (or viewer) count, live-stream watch time and short video watch time.
@@
To address the aforementioned issues, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm, SL-MGAC. Given the significant variation in users’ interest in live-stream content, which introduces high variance at both the feature and feedback levels, we incorporate separate multi-group state decomposition (MG-SD) modules within both the actor and critic networks. Additionally, we combine multi-task supervised reward learning with traditional critic learning to not only restrict temporal difference (TD) error accumulation during model training but also improve the accuracy of Q-value estimation. Through experiments, we compare the SL-MGAC method with competitive baselines using offline policy evaluation (OPE) (Uehara et al., 2022) and online A/B testing to demonstrate its effectiveness and stability.
@@
  • We propose new variance reduction (Greensmith et al., 2004; Kumar et al., 2019; Anschel et al., 2017; Mao et al., 2018; Romoff et al., 2018) techniques, including multi-group state decomposition, distribution discretization for reward learning, reward normalization, and Q-value normalization.
@@
  • We successfully deploy the SL-MGAC model in a challenging short video and live-stream mixed recommendation scenario for Kwai, a short video app with over 100 million users.
@@ 2. Problem Formulation
As shown in Fig. 1, in the short video & live-stream mixed recommendation system, the live-stream recommendation system is treated as an agent that interacts with diverse users and receives user feedback over time. The optimal live-stream injection control problem is modeled as an infinite request-level MDP to maximize the cumulative reward. Formally, we define the live-stream injection MDP as a tuple of five elements $\left(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\right)$:
  • State space $\mathcal{S}$: This is the set of user interaction states $s$, which includes user static features (e.g., user ID, location, gender, country), user history features (e.g., live-stream watch history, short video watch history), and item features (e.g., live-stream ID, author ID, author gender, online viewer count). We limit the length of users’ live-stream and short video history lists by timestamp order, keeping only the top 50 items.
  • Action space $\mathcal{A}$: The action $a\in\mathcal{A}$ represents the decision of whether to inject a recommended live-stream in response to a user’s request. We define action $a$ as a binary variable, where $a=1$ means injecting a live-stream, and $a=0$ means not injecting it.
@@
  • Discount Factor $\gamma$: The discount factor $\gamma\in[0,1]$ is used in the computation of cumulative reward. Typically, we set $\gamma<1$ during model training.
@@
Overall, the optimization objective of aforementioned optimal live-stream injection control problem is shown as follows:
(1) $\max_{\pi}\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$
where $\pi(a_{t}|s_{t})$ is the live-stream injection policy to be optimized.
2.1. Reward Function Design
In our mixed short video & live-stream recommendation scenario, we design a reward function based on the time gain of live-stream watch time relative to the average short video watch time in a request:
(2) $r(s_{t},a_{t})=\begin{cases}y_{l}-\frac{1}{B}y_{v},\quad&if\;a_{t}=1\\ \beta&otherwise\\ \end{cases}$
where $y_{l}$ is the live-stream watch time, $y_{v}$ is the total video watch time, $B$ is the number of videos in a request, and $\beta$ is the hyper-parameter in Eq. 2. Note that the recommended list in a request always contains $B$ videos and at most one live-stream. Furthermore, the reward defined above is similar to that of the RCPO (Tessler et al., 2018) algorithm. From the perspective of constrained reinforcement learning, the term $\frac{1}{B}y_{v}$ can be viewed as an implicit cost constraint. In practice, the hyper-parameter $\beta$ is introduced to carefully control the final live-stream injection ratio.
Since the distribution of live-stream (short video) watch time changes drastically over time, as shown in Fig. 2b, it is challenging for an RL agent to optimize the cumulative reward in Eq. 2. To address this, we propose a reward normalization technique as follows:
(3) $r(s_{t},a_{t})=\begin{cases}sigmoid\left(y_{l}-\frac{1}{B}y_{v}\right),\quad&if\;a_{t}=1\\ \beta\in[0,1]&otherwise\\ \end{cases}$
which helps stabilize the network gradient, thereby reducing the risk of model divergence.
@@ 3. Proposed Framework
To address the intrinsic divergence and instability issues of reinforcement learning and successfully deploy the RL agent in our mixed short video & live-stream recommendation system, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC), as shown in Fig. 3. SL-MGAC incorporates a supervised learning-enhanced actor-critic model with a shared user & live-stream feature extraction module, along with independent Multi-Group State Decomposition (MG-SD) modules. For clarity, only one critic network is shown in Fig. 3, while the remaining three critic networks, which use the same architecture, are omitted.
3.1. User & Live-stream Feature Extraction Module
The user & live-stream feature extraction module aims to generate non-linear embedding representations for the state $s_{t}$. First, we use a unified embedding layer to map the user’s static features, live-stream features, historical live-stream list features and short video list features into low-dimensional dense representations. We denote $\mathbf{v}_{U}$, $\mathbf{v}_{L}$, $\left\{\mathbf{e}_{1},\cdots,\mathbf{e}_{M_{l}}\right\}$, $\left\{\mathbf{e}^{\prime}_{1},\cdots,\mathbf{e}^{\prime}_{M_{v}}\right\}$ as the corresponding embedding vectors or sets of embedding vectors, respectively. We then define $\mathbf{h}_{a}=[\mathbf{v}_{U},\mathbf{v}_{L}]$ as the concatenation of $\mathbf{v}_{U}$ and $\mathbf{v}_{L}$.
To aggregate historical live-stream (short video) representations, we introduce the target attention mechanism (Vaswani, 2017; Zhou et al., 2018), which is defined as follows:
@@ 3.1. User & Live-stream Feature Extraction Module
After aggregating the historical live-stream and short video features via two separate attention networks, we concatenate all the above embedding vectors to form $\mathbf{h}_{s}=[\mathbf{v}_{U},\mathbf{v}_{L},\mathbf{h}_{live},\mathbf{h}_{video}]$. We then use a shared multi-layer perceptron (MLP) for both the actor and critic networks to obtain the latent representation of state $s_{t}$, i.e. $\mathbf{h}^{\prime}_{s}=f_{MLP}(\mathbf{h}_{s})$. To ensure training stability, we stop the gradient flow through $\mathbf{h}^{\prime}_{s}$ for the actor network and only use the more complex critic networks to optimize $\mathbf{h}^{\prime}_{s}$, as we find that sharing the feature extraction module and the $f_{MLP}$ network causes interference between the policy and critic gradients.
@@ 3.2. Multi-Group State Decomposition Module
Since live-stream content is not equally appealing to all users on the short video app Kwai, it is crucial to inject live-streams selectively into the short video feed. Otherwise, excessive live-stream exposures may disrupt a user’s interest in short videos, leading to a decrease in overall app usage duration. In practice, user interaction data for live-streams is sparser and noisier compared to that for short videos, making it challenging for RL algorithms to learn an optimal live-stream injection policy for each individual user.
Therefore, prior information about diverse user groups is essential for enhancing the decision-making accuracy of RL models. A natural approach to distinguish between different user behaviors is to partition users into distinct groups based on their historical activity. Specifically, we categorize users into $K$ disjoint groups according to their live-stream activity level, which is determined by the cumulative live-stream watch time over the past 3 weeks. Users with higher historical watch time are assigned to groups with higher activity levels.
Formally, we present the following theorem, which demonstrates that the proposed multi-group state decomposition can be viewed as a type of variance reduction technique for real-world RL applications.
Theorem 3.1.
Let state space $\mathcal{S}=\bigcup_{i=1}^{K}\mathcal{S}_{i},\forall i\neq j\;\mathcal{S}_{i}\cap\mathcal{S}_{j}=\varnothing$, state $s\in\mathcal{S}$; $r$ be the real reward of state $s$ and action $a$, $\hat{r}$ be the predicted value of $r$ over $\mathcal{S}$, $\hat{r}_{i}$ be the predicted reward of $r$ over $\mathcal{S}_{i}$; then a neural network $f$ with MG-SD module has a lower prediction variance than the neural network $f^{\prime}$ without MG-SD module.
Please refer to Appendix. A for detailed proof of Theorem 3.1. We demonstrate that the MG-SD module eliminates between-group variance, thereby reducing the total variance, which implies that the multi-group state decomposition module contributes to a more stable RL model. Empirically, we also show the effectiveness of this module in enhancing RL model stability in subsequent experiments. We find that SL-MGAC without the MG-SD module tends to be more unstable during training, with a larger variance in its Q-values.
@@ 3.3.1. Layer Normalization
A recent work on RL divergence and instability (Yue et al., 2024) shows that offline critic learning exhibits a self-excitation pattern. Specifically, Q-value iteration can inadvertently cause the target Q value $Q(s_{t+1},a^{*})$ to increase even more than the increment of $Q(s_{t},a_{t})$, which amplifies the TD-error and traps the critic learning process in a positive feedback loop until model collapse. Therefore, normalization techniques, such as Layer Normalization (Ba, 2016) can be utilized to alleviate the divergence and instability problems.
From the proof in the Appendix D of (Yue et al., 2024), we know that for any input $\mathbf{x}$ and any direction $\mathbf{v}$, if a network applies Layer Normalization to the input, then we have $k_{\rm NTK}(\mathbf{x},\mathbf{x}+\lambda\mathbf{v})\leq C$, where $\lambda>0$, $C$ is a constant and $k_{\rm NTK}$ is the Neural Tangent Kernel (NTK) (Jacot et al., 2018). This theoretically indicates that a network with Layer Normalization is less sensitive to input variations and can maintain stable gradients despite perturbations during model training. Hence, we apply Layer Normalization to the inputs of both the actor and critic networks, as shown in Fig. 3.

### v2

# Supervised Learning-enhanced Multi-Group Actor Critic for Live Stream Allocation in Feed

@@ Abstract.
Reinforcement Learning (RL) has been widely applied in recommendation systems to capture long-term user engagement, thus improving dwelling time and improving user retention. In the context of a short video & live stream mixed recommendation scenario, the live stream recommendation system (RS) decides whether to inject at most one live stream into the video feed for each user request. To maximize long-term user engagement, it is crucial to determine an optimal live stream injection policy for accurate live stream allocation. However, traditional RL algorithms often face divergence and instability problems, and these issues may cause too many live stream allocation, which interrupts user’s short video interest and leads to a decrease in the user’s app usage duration. To address these challenges, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm (SL-MGAC). Specifically, we introduce a supervised learning-enhanced actor-critic framework that incorporates variance reduction techniques, where multi-task reward learning helps restrict bootstrapping error accumulation during critic learning. Additionally, we design a multi-group state decomposition module for both actor and critic networks to reduce prediction variance and improve model stability. We also propose a novel reward function to prevent overly greedy live stream allocation. Empirically, we evaluate the SL-MGAC algorithm using offline policy evaluation (OPE) and online A/B testing. Experimental results demonstrate that the proposed method not only outperforms baseline methods but also exhibits enhanced stability in online recommendation scenarios.
@@
Figure 1. Structure of a short video & live stream mixed recommendation system(RS). The decision making of SL-MGAC takes place in the final stage of live stream RS.
@@ 1. Introduction
Deep reinforcement learning has shown great potential in learning policies to maximize long-term rewards in various research domains, such as computer vision, robotics, natural language processing, gaming, and recommendation systems (Afsar et al., 2022; Li, 2017). In the context of recommendation systems, reinforcement learning is applied to optimize long-term user engagement (Zou et al., 2019) and improve user retention (Cai et al., 2023). Numerous RL applications have been proposed for real-world recommendation systems, including slate recommendation (Ie et al., 2019; Deffayet et al., 2023; Liu et al., 2023), personalized search ranking (Miao et al., 2021), and advertisement allocation (Liao et al., 2022).
@@
We consider a challenging sequential decision making scenario in a short video & live stream blended recommendation system, as shown in Fig. 1. The system consists of three components: a live streaming recommendation system, a short video recommendation system, and a blending server. For each user request with timestamp $t$, the live stream recommendation system decides whether to inject the recommended live stream into the video feed, while the short video recommendation system suggests $B$ (with $B<10$) videos. The blending server then mixes and rearranges the single live stream with the $B$ videos to form the final recommendation list. Our objective is to find a personalized optimal live stream injection policy that not only maximizes user long-term engagement with live streams, but also meets the constraint of not causing a negative impact on the app usage duration, which can be modeled as an infinite request-level Constrained Markov Decision Process (CMDP) (Altman, 2021).
@@
However, the intrinsic issues of divergence and instability associated with traditional reinforcement learning algorithms (RL) (François-Lavet et al., 2015; Fujimoto et al., 2018; Kumar et al., 2019; Dasagi et al., 2019) are significantly exacerbated in this challenging scenario. A primary possible reason for this is the drastic fluctuations in both the live stream supply scale and user interaction behaviors over time, as shown in Fig. 7 in Appendix B. RL models often struggle to learn effective policies from data that exhibit such high variance across both time scales and user scales. In practice, we observe that RL models frequently encounter issues of policy deterioration or model collapse (Dohare et al., 2023).
Furthermore, like advertisement, live stream is not welcomed by every user. Allocating too many live streams for a user will significantly interrupt the user’s short video interest and lead to a decrease in the user’s app usage duration, which will eventually affect the user engagement and user retention. Therefore, we cannot merely focus on the user feedback for a single request, but rather should optimize the long-term user experience.
@@
To address the aforementioned issues, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm, SL-MGAC. Given the significant variation in users’ interest in live stream content, which introduces high variance at both the feature and feedback levels, we incorporate separate multi-group state decomposition (MG-SD) modules within both the actor and critic networks. Additionally, we combine multi-task supervised reward learning with traditional critic learning to not only restrict temporal difference (TD) error accumulation during model training but also improve the accuracy of Q-value estimation. Through experiments, we compare the SL-MGAC method with competitive baselines using offline policy evaluation (OPE) (Uehara et al., 2022) and online A/B testing to demonstrate its effectiveness and stability.
@@
  • We propose new variance reduction (Greensmith et al., 2004; Kumar et al., 2019; Anschel et al., 2017; Mao et al., 2018; Romoff et al., 2018) techniques, including multi-group state decomposition, distribution discretization for reward learning, reward normalization, and Q value normalization.
@@
  • We successfully deploy the SL-MGAC model in a challenging short video and live stream mixed recommendation scenario for Kwai, a short video app with over 100 million users.
@@ 2. Problem Formulation
As shown in Fig. 1, in the short video & live stream mixed recommendation system, the live stream recommendation system is treated as an agent that interacts with various users and receives user feedback over time. However, an agent considering instant user feedback of a single request may allocate live streams greedily in a short video feed and eventually affect user retention. Therefore, the optimal live stream injection control problem is modeled as an infinite request level CMDP to maximize the cumulative live stream reward and satisfy the platform-level constraint, for example, live stream allocation cannot reduce the total app usage duration.
Formally, we define live stream injection CMDP as a tuple of six elements $\left(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\mathcal{C}\right)$:
  • State space $\mathcal{S}$: This is the set of user interaction states $s$, which includes user static features (e.g., user ID, location, gender, country), user history features (e.g., live stream watch history, short video watch history) and item features (e.g., live stream ID, author ID, author gender, online viewer count). We limit the length of user live stream and short video history lists by timestamp order, keeping only the top 50 items.
  • Action space $\mathcal{A}$: The action $a\in\mathcal{A}$ represents the decision of whether to inject a recommended live stream in response to a user’s request. We define the action $a$ as a binary variable, where $a=1$ means injecting a live stream and $a=0$ means not injecting it.
@@
  • Discount Factor $\gamma$: The discount factor $\gamma\in[0,1]$ is used in the calculation of the cumulative reward. Typically, we set $\gamma<1$ during model training.
  • Constraint $\mathcal{C}$: The platform-level constraint is that the live stream watch time should be longer than the average video watch time per request as much as possible. This prevents a negative impact on app usage duration, which could happen if the live stream watch time is shorter. By meeting this constraint, we ensure that live streams boost users’ long-term app engagement instead of reducing their time spent on it.
@@
Overall, the optimization objective of aforementioned optimal live stream injection control problem is shown as follows:
(1) $\max_{\pi}J_{R}(\pi)$
$s.t.\quad J_{C}(\pi)\leq\epsilon$
$J_{R}(\pi)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$
$J_{C}(\pi)=\frac{1}{B}y_{v}-y_{l}$
where $\pi(a_{t}|s_{t})$ is the live stream injection policy to be optimized, $y_{l}$ is the live stream watch time, $y_{v}$ is the total video watch time, and $B$ is the number of videos in a request.
A common method to solve the above CMDP is to transform the problem into a min-max optimization by introducing the Lagrange multiplier $\lambda$:
(2) $(\pi^{*},\lambda^{*})=\arg\min_{\lambda\geq 0}\max_{\pi}J_{R}(\pi)-\lambda(J_{C}(\pi)-\epsilon)$
Inspired by the RCPO (Tessler et al., 2018) algorithm, the primal-dual optimization problem can be transformed into an equivalent unconstrained problem with penalized reward function:
(3) $\hat{r}(\lambda,s_{t},a_{t})$ $=r(s_{t},a_{t})-\lambda c(s_{t},a_{t})$
$=y_{l}-\lambda(\frac{1}{B}y_{v}-y_{l})$
$=(1+\lambda)y_{l}-\frac{\lambda}{B}y_{v}$
In practice, we solve the unconstrained MDP problem and the reward function is simplified as follows:
(4) $r(s_{t},a_{t})=y_{l}-\frac{\lambda}{B}y_{v}$
where $\lambda$ is a hyper-parameter to carefully control the final live stream injection ratio. From the perspective of constrained reinforcement learning, the term $\frac{\lambda}{B}y_{v}$ can be viewed as an implicit cost constraint.
@@ 3. Proposed Framework
To address the intrinsic divergence and instability issues of reinforcement learning and successfully deploy the RL agent in our mixed short video & live stream recommendation system, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC), as shown in Fig. 2. SL-MGAC incorporates a supervised learning-enhanced actor-critic model with a shared user & live stream feature extraction module, along with independent Multi-Group State Decomposition (MG-SD) modules. For clarity, only one critic network is shown in Fig. 2, while the remaining three critic networks, which use the same architecture, are omitted.
3.1. User & live stream Feature Extraction Module
The user & live stream feature extraction module aims to generate non-linear embedding representations for the state $s_{t}$. First, we use a unified embedding layer to map the user’s static features, live stream features, historical live stream list features, and short video-list features into low-dimensional dense representations. We denote $\mathbf{v}_{U}$, $\mathbf{v}_{L}$, $\left\{\mathbf{e}_{1},\cdots,\mathbf{e}_{M_{l}}\right\}$, $\left\{\mathbf{e}^{\prime}_{1},\cdots,\mathbf{e}^{\prime}_{M_{v}}\right\}$ as the corresponding embedding vectors or sets of embedding vectors, respectively. We then define $\mathbf{h}_{a}=[\mathbf{v}_{U},\mathbf{v}_{L}]$ as the concatenation of $\mathbf{v}_{U}$ and $\mathbf{v}_{L}$.
To aggregate historical live stream (short video) representations, we introduce the target attention mechanism (Vaswani, 2017; Zhou et al., 2018), which is defined as follows:
@@ 3.1. User & live stream Feature Extraction Module
After aggregating the historical live stream and short video features via two separate attention networks, we concatenate all of the above embedding vectors to form $\mathbf{h}_{s}=[\mathbf{v}_{U},\mathbf{v}_{L},\mathbf{h}_{live},\mathbf{h}_{video}]$. We then use a shared multi-layer perceptron (MLP) for both the actor and critic networks to obtain the latent representation of state $s_{t}$, i.e. $\mathbf{h}^{\prime}_{s}=f_{MLP}(\mathbf{h}_{s})$. To ensure training stability, we stop the gradient flow through $\mathbf{h}^{\prime}_{s}$ for the actor network and only use the more complex critic networks to optimize $\mathbf{h}^{\prime}_{s}$, as we find that sharing the feature extraction module and the $f_{MLP}$ network causes interference between the policy and critic gradients.
@@ 3.2. Multi-Group State Decomposition Module
Since live stream content is not equally appealing to all users of the short video app Kwai, it is crucial to inject live streams selectively into the short video feed. Otherwise, excessive live stream exposures may disrupt a user’s interest in short videos, leading to a decrease in overall app usage duration. In practice, user interaction data for live streams are sparser and noisier compared to that for short videos, making it challenging for RL algorithms to learn an optimal live stream injection policy for each individual user. Directly learning a policy from sparse and noisy user feedback is often infeasible, since it will inadvertently inject excessive live streams in the short video feed and highly affected the user engagement.
We find that prior information on diverse user groups is essential to improve the accuracy of the decision making of RL models in our scenario, which provides additional user preference information to the policy. A natural approach to distinguish between different user behaviors is to partition users into distinct groups based on their historical activity. Specifically, we categorize users into $K$ disjoint groups according to their activity level of the live stream, which is determined by the cumulative watch time of the live stream during the past 3 weeks. Users with a higher historical watch time are assigned to groups with higher activity levels.
Empirically, we show the effectiveness of this module in enhancing the stability of the RL model in subsequent experiments. We find that SL-MGAC without the MG-SD module tends to be more unstable during training, with a larger variance in its Q-values.
@@ 3.3.1. Layer Normalization
A recent work on RL divergence and instability (Yue et al., 2024) shows that offline critic learning exhibits a self-excitation pattern. Specifically, Q-value iteration can inadvertently cause the target Q value $Q(s_{t+1},a^{*})$ to increase even more than the increment of $Q(s_{t},a_{t})$, which amplifies the TD-error and traps the critic learning process in a positive feedback loop until model collapse. Therefore, normalization techniques, such as Layer Normalization (Ba, 2016) can be utilized to alleviate divergence and instability problems.
From the proof in Appendix D of (Yue et al., 2024), we know that for any input $\mathbf{x}$ and any direction $\mathbf{v}$, if a network applies Layer Normalization to the input, then we have $k_{\rm NTK}(\mathbf{x},\mathbf{x}+\lambda\mathbf{v})\leq C$, where $\lambda>0$, $C$ is a constant, and $k_{\rm NTK}$ is the Neural Tangent Kernel (NTK) (Jacot et al., 2018). This theoretically indicates that a network with Layer Normalization is less sensitive to input variations and can maintain stable gradients despite perturbations during model training. Hence, we apply Layer Normalization to the inputs of both the actor and critic networks, as shown in Fig. 2.